# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 3 Learning Multi-Index Models with Agnostic Noise
Fix $i\in\mathcal{Y}$ and consider the boolean functions $f_{i}(\mathbf{x})\coloneqq\mathds{1}(f(\mathbf{x})=i)$, $g_{i}(\mathbf{x})\coloneqq\mathds{1}(g(\mathbf{x}^{V})=i)$.
By $T_{\rho}$ we denote the Gaussian noise operator defined in Definition A.3. By A.4, we have that $\left\|f_{i}-g_{i}\right\|_{L_{1}}\geq\left\|T_{\rho}f_{i}-T_{\rho}g_{i}\right\|_{L_{1}}$. Hence, from the triangle inequality we also have that $\left\|f_{i}-g_{i}\right\|_{L_{1}}+\left\|T_{\rho}f_{i}-f_{i}\right\|_{L_{1}}\geq\left\|f_{i}-T_{\rho}g_{i}\right\|_{L_{1}}$. Furthermore, by A.5, we have that $\left\|T_{\rho}f_{i}-f_{i}\right\|_{L_{1}}\leq\sqrt{\pi}\Gamma(f_{i})\rho$. Therefore, we have that
@@
Moreover, note that for any continuously differentiable function $F:\mathbb{R}^{d}\mapsto\mathbb{R}$, it holds that $F(\mathbf{a})-F(\mathbf{b})=\int_{0}^{1}\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\cdot(\mathbf{a}-\mathbf{b})\mathrm{d}t$ and from the Cauchy-Schwarz inequality, we have that $|F(\mathbf{a})-F(\mathbf{b})|\leq\max_{t\in(0,1)}\|\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\|\|\mathbf{a}-\mathbf{b}\|$. By taking $F=T_{\rho}g_{i}$ and using A.4, we get that
$|T_{\rho}g_{i}(\mathbf{x})-T_{\rho}g_{i}(\mathbf{y})|\leq\frac{\|\mathbf{x}-\mathbf{y}\|}{\rho}\;,$
for any $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$. For each $S\in\mathcal{S}$, pick a point $\mathbf{x}_{S}\in S$. By averaging over the partition, we have that
@@
Furthermore, observe that by the linearity of expectation and the fact that $\sum_{i\in\mathcal{Y}}g_{i}(\mathbf{x})=1$ we have that $\sum_{i\in\mathcal{Y}}T_{\rho}g_{i}(\mathbf{x})=1$ for $\mathbf{x}\in\mathbb{R}^{d}$. Moreover, $T_{\rho}g_{i}(\mathbf{x})\in[0,1]$ for all $\mathbf{x}\in\mathbb{R}^{d},i\in\mathcal{Y}$. Therefore, we can consider the multiclass function $G:\mathbb{R}^{d}\to\mathcal{Y}$ such that $G(\mathbf{x})$ is a random label sampled from the probability distribution induced by $T_{\rho}g_{i}(\mathbf{x}_{S})$, where $S\in\mathcal{S}$ such that $\mathbf{x}\in S$. Note that for every probability vector $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ and label $z\in\mathcal{Y}$ we have that
@@
Noting that $\sum_{i\in\mathcal{Y}}\|f_{i}-g_{i}\|=2\tau$ set $\rho=\epsilon^{\prime}/(C{\Gamma(f)})$ and $\eta={\epsilon^{\prime}\rho}/{(C|\mathcal{Y}|\sqrt{k}})$, for a sufficiently large constant $C>0$. With these choices, we obtain $\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq f(\mathbf{x}),\mathbf{x}\in A]\leq\tau+\epsilon^{\prime}/2$. By the triangle inequality we have that $\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq y,\mathbf{x}\in A]]\leq\tau+\epsilon+\epsilon^{\prime}/2$. Since $h_{\mathcal{S}}$ chooses the label that minimizes the error at each $S\in\mathcal{S}$, we have that $\operatorname*{\mathbf{Pr}}[h_{\mathcal{S}}(\mathbf{x})\neq y,\mathbf{x}\in A]\leq\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq y,\mathbf{x}\in A]]$. Finally, considering that $\operatorname*{\mathbf{Pr}}[\mathbf{x}\not\in A]\leq\eta$ completes the proof of Lemma 3.4.
In the following lemma, we show that, due to the bound on the Gaussian surface area of $f$, if there exist level sets of $f$ for which the conditional moments on a subspace $V$ are large with non-trivial probability over $V$, then the same property holds on sufficiently small cubes. That is, when $V$ is partitioned into sufficiently thin cubes, the conditional moments of the level sets of $f$ on these cubes will be large with non-trivial probability over the choice of cube. The intuition behind the proof is that, since the function is smooth (by the GSA bound), re-randomizing over sufficiently small cubes barely affects its output. However, by re-randomizing over cubes, the moments of the level-sets become independent of which point in the cube we condition on. As a result, if we observe distinguishing moments when conditioning on a point in the cube, then these moments persist when conditioning on the entire cube.
@@
Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and let $\mathcal{F}$ be a class of $(m,\epsilon,\delta,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2. Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$, with $k\geq 1$, and let $\mathcal{S}$ be an $\eta$-approximating partition of $V$ according to Definition 2.2, with $\eta=\epsilon^{\prime 3}/(C\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2})$, for $C>0$ a sufficiently large universal constant.
If for a function $f\in\mathcal{F}$ with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\epsilon+\epsilon^{\prime}$ there exist no function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$, then there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\delta$ such that for each $S\in\mathcal{T}$ there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma\|p(\mathbf{x}^{U})\|_{2}$, where $U=(V+W)\cap V^{\perp}$.
@@
Note that, since $\mathcal{F}$ satisfies Definition 3.2, with probability at least $\delta$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, independent of $\mathbf{x}$, there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that
@@
In order to prove the statement, we need to show that the above condition can be extended from a single point to encompass an entire cube $S$. To achieve this, we introduce a random variable $y^{\prime}$, which is $\epsilon$-close to $f$, such that, when conditioned on any $S\subset\mathcal{S}$, $y^{\prime}$ is distributed uniformly within $S$ along the directions defined by $V$, while the other directions remain consistent with those of $y$.
@@
In the following claim we will prove that $f(\mathbf{x})$ and $f(\mathbf{x}^{\prime})$ are approximately the same.
@@
It holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]=O(\epsilon^{\prime})$.
@@
Define the boolean function $f_{i}(\mathbf{x})\coloneqq\mathds{1}(f(\mathbf{x})=i)$ and denote by $g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})\coloneqq T_{\rho}f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})$, where $T_{\rho}$ denotes the Gaussian noise operator defined in Definition A.3. Fix $\mathbf{x}_{0}\in\mathbb{R}^{d},S\in\mathcal{S}$ we first show that $\lvert g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})-g_{\mathbf{x}_{0}}^{(i)}(\mathbf{z})\rvert$ is small for $\mathbf{x}^{V},\mathbf{z}^{V}\in S$.
Note that for any continuously differentiable function $F:V\mapsto\mathbb{R},\mathbf{a},\mathbf{b}\in V$, it holds that $F(\mathbf{a})-F(\mathbf{b})=\int_{0}^{1}\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\cdot(\mathbf{a}-\mathbf{b})\mathrm{d}t$ and from Cauchy-Schwarz inequality, we have that $|F(\mathbf{a})-F(\mathbf{b})|\leq\max_{t\in(0,1)}\|\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\|\|\mathbf{a}-\mathbf{b}\|$. By taking $F(\mathbf{x})=g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})$ and using A.4, we get that
@@
for any $\mathbf{x}^{V},\mathbf{\mathbf{z}}^{V}\in S$.
Consequently by the triangle inequality and Claim A.7 we have
$\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{E}}_{\mathbf{x}_{0},\mathbf{x}^{V}}[\lvert(f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x}))\mathds{1}(\mathbf{x}\in S)\rvert]$
@@
$\leq\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{E}}_{\mathbf{x}_{0},\mathbf{x}^{V}}[\lvert(f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})-g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x}_{S}))\mathds{1}(\mathbf{x}\in S)\rvert]+\eta/\rho\lesssim\sqrt{\rho}\Gamma(f_{i})+\eta\sqrt{k}/\rho\;,$
@@
for any collection of points $\mathbf{x}_{S}\in S$. Furthermore, observe that by the linearity of expectation and the fact that $\sum_{i\in\mathcal{Y}}f_{i}(\mathbf{x})=1$ we have that $\sum_{i\in\mathcal{Y}}g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})=1$ and $g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})\in[0,1]$ for all $\mathbf{x},\mathbf{x}_{0}\in\mathbb{R}^{d},i\in\mathcal{Y}$. Therefore, we can consider the multiclass function $G:\mathbb{R}^{d}\to\mathcal{Y}$ such that $G(\mathbf{x})$ is a random label sampled from the probability distribution induced by $g_{\mathbf{x}}^{(i)}(\mathbf{x}_{S})$, where $S\in\mathcal{S}$ such that $\mathbf{x}\in S$. Note that for every probability vector $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ and label $z\in\mathcal{Y}$ we have that
Therefore, we have that the $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq G(\mathbf{x})]\lesssim\sqrt{\rho}\Gamma(f)+\eta\sqrt{k}\lvert\mathcal{Y}\rvert/\rho\lesssim\epsilon^{\prime}$.
@@
Notice that since $\mathbf{x}^{\prime}$ is distributed as $\mathcal{N}(\mathbf{0},\mathbf{I})$, we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x}^{\prime})]=O(\epsilon^{\prime})$. However, as $G(\mathbf{x})$ is independent $\mathbf{x}^{V}$, we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x}^{\prime})]=\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x})]=O(\epsilon^{\prime})$. Therefore, by the triangle inequality we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]=O(\epsilon^{\prime})$, which completes the proof of Claim 3.6.∎
@@
Define $y^{\prime}$ such that or all $\mathbf{z}\in\mathbb{R}^{d}$, the distribution of $y^{\prime}$ given $\mathbf{x}^{\prime}=\mathbf{z}$ is the same as the distribution of $y$ given $\mathbf{x}=\mathbf{z}$. By the definition of $y^{\prime}$, it follows that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}^{\prime},y^{\prime}}[f(\mathbf{x}^{\prime})\neq y^{\prime}]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}^{\prime}}\left[\operatorname*{\mathbf{E}}_{y^{\prime}}[f(\mathbf{x}^{\prime})\neq y^{\prime}\mid\mathbf{x}^{\prime}]\right]\leq\epsilon$. Therefore, it holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y^{\prime}]\leq\epsilon+O(\epsilon^{\prime})$.
@@
Therefore, with probability at least $\delta$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ independent of $\mathbf{x}$, there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that
@@
However, since the moments of $y^{\prime}$ are identical when conditioned on each $\mathbf{x}_{0}^{V}$ that belongs to the same region in $S$, we deduce that if the moment condition holds for some polynomial at a particular point $\mathbf{x}_{0}^{V}$, then it must hold for the entire cube $S$, such that $\mathbf{x}_{0}^{V}\in S$. Therefore, there exists a subset $\mathcal{T}\subseteq\mathcal{S}$ of the partition with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\delta$ satisfying the following: for all $S\in\mathcal{T}$ there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]>\sigma\|p(\mathbf{x}^{U})\|_{2}\;.$
@@
The remaining part of the section is structured as follows: In Section 3.1, we provide the analysis of Algorithm 4 with Proposition 3.7. In Section 3.2, we analyze Algorithm 3 providing the full proof of Theorem 3.3, which is the main result of this section.

### v2

@@ 3 Learning Multi-Index Models with Agnostic Noise
Fix $i\in\mathcal{Y}$ and $\rho\in(0,1)$ and consider the boolean functions $f_{i}(\mathbf{x})\coloneqq\mathds{1}(f(\mathbf{x})=i)$, $g_{i}(\mathbf{x})\coloneqq\mathds{1}(g(\mathbf{x}^{V})=i)$. We denote the Gaussian noise operator by $T_{\rho}$, as defined in Definition A.3. By A.4, we have that $\left\|f_{i}-g_{i}\right\|_{L_{1}}\geq\left\|T_{\rho}f_{i}-T_{\rho}g_{i}\right\|_{L_{1}}$. Hence, from the triangle inequality we also have that $\left\|f_{i}-g_{i}\right\|_{L_{1}}+\left\|T_{\rho}f_{i}-f_{i}\right\|_{L_{1}}\geq\left\|f_{i}-T_{\rho}g_{i}\right\|_{L_{1}}$. Furthermore, by the Ledoux-Pisier inequality (A.5), we have that $\left\|T_{\rho}f_{i}-f_{i}\right\|_{L_{1}}\leq\sqrt{\pi}\Gamma(f_{i})\rho$. Therefore, we have that
@@
Moreover, note that for any continuously differentiable function $F:\mathbb{R}^{d}\mapsto\mathbb{R}$, it holds that $F(\mathbf{a})-F(\mathbf{b})=\int_{0}^{1}\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\cdot(\mathbf{a}-\mathbf{b})\mathrm{d}t$. By the Cauchy-Schwarz inequality, we have that $|F(\mathbf{a})-F(\mathbf{b})|\leq\max_{t\in(0,1)}\|\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\|\|\mathbf{a}-\mathbf{b}\|$. By taking $F=T_{\rho}g_{i}$ and using A.4, we have that for any $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, it holds
$|T_{\rho}g_{i}(\mathbf{x})-T_{\rho}g_{i}(\mathbf{y})|\leq\frac{\|\mathbf{x}-\mathbf{y}\|}{\rho}\;.$
For each $S\in\mathcal{S}$, pick a point $\mathbf{x}_{S}\in S$. Then for any $\mathbf{x}\in S$, the distance between $\mathbf{x}$ and $\mathbf{x}_{S}$ is at most the diameter of $S$. In our setting, since the partition is constructed along a $k$-dimensional subspace $V$ with edge length roughly $\eta$ (i.e., each coordinate is in an interval of length $\eta$), the distance between any $\mathbf{x}\in S$ and $\mathbf{x}_{S}$ is at most $\|\mathbf{x}-\mathbf{x}_{S}\|\leq\sqrt{k}\eta.$ Thus, for every $\mathbf{x}\in S$
$|T_{\rho}g_{i}(\mathbf{x})-T_{\rho}g_{i}(\mathbf{x}_{S})|\leq\frac{\|\mathbf{x}-\mathbf{x}_{S}\|}{\rho}\leq\frac{\sqrt{k}\cdot\eta}{\rho}.$
By averaging over the partition $\mathcal{S}$, we have that
@@
Furthermore, by linearity of expectation and the fact that $\sum_{i\in\mathcal{Y}}g_{i}(\mathbf{x})=1$ we have that $\sum_{i\in\mathcal{Y}}T_{\rho}g_{i}(\mathbf{x})=1$ for $\mathbf{x}\in\mathbb{R}^{d}$. Moreover, $T_{\rho}g_{i}(\mathbf{x})\in[0,1]$ for all $\mathbf{x}\in\mathbb{R}^{d},i\in\mathcal{Y}$.
Now, we define a multiclass function $G:\mathbb{R}^{d}\to\mathcal{Y}$ in the following way: for each point $\mathbf{x}\in\mathbb{R}^{d}$, $G(\mathbf{x})$ is a random label chosen according to the probability distribution induced by $T_{\rho}g_{i}(\mathbf{x}_{S})$, where $S\in\mathcal{S}$ is the region such that $\mathbf{x}\in S$. Furthermore, for every probability vector $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ (where each element of $\mathbf{p}$ represents the probability of the corresponding label) and the unit vector $\mathbf{e}_{z}$, which represents the one-hot encoding of the label $z\in\mathcal{Y}$. We have that
@@
Noting that $\sum_{i\in\mathcal{Y}}\|f_{i}-g_{i}\|=2\tau$, set $\rho=\epsilon^{\prime}/(C{\Gamma(f)})$ and $\eta={\epsilon^{\prime}\rho}/{(C|\mathcal{Y}|\sqrt{k}})$, for a sufficiently large constant $C>0$. With these choices, we obtain $\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq f(\mathbf{x}),\mathbf{x}\in A]\leq\tau+\epsilon^{\prime}/2$. By the triangle inequality we have that $\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq y,\mathbf{x}\in A]]\leq\tau+\epsilon+\epsilon^{\prime}/2$. Since $h_{\mathcal{S}}$ chooses the label that minimizes the error at each $S\in\mathcal{S}$, we have that $\operatorname*{\mathbf{Pr}}[h_{\mathcal{S}}(\mathbf{x})\neq y,\mathbf{x}\in A]\leq\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq y,\mathbf{x}\in A]$. Finally, considering that $\operatorname*{\mathbf{Pr}}[\mathbf{x}\not\in A]\leq\eta$ (recall that $\eta\leq\epsilon^{\prime 2}$) completes the proof of Lemma 3.4. ∎
In the following lemma, we show that, due to the bound on the Gaussian surface area of $f$, if there is a non-negligible probability that the conditional moments of $f$’s level sets along a subspace $V$ are large, the same property holds when $V$ is partitioned into sufficiently small cubes. That is, if we partition $V$ into thin cubes, then with non-trivial probability the conditional moments computed over each cube remain large. The intuition behind the proof is that, since the function is smooth (by the GSA bound), re-randomizing over sufficiently small cubes barely affects its output. However, by re-randomizing over cubes, the moments of the levelsets become independent of which point in the cube we condition on. As a result, if we observe distinguishing moments when conditioning on a point in the cube, then these moments persist when conditioning on the entire cube.
@@
Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and let $\mathcal{F}$ be a class of $(m,\zeta+O(\epsilon),\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2. Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$, with $k\geq 1$, and let $\mathcal{S}$ be an $\eta$-approximating partition of $V$ according to Definition 2.2, with $\eta=\epsilon^{3}/(C\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2})$, for $C>0$ a sufficiently large universal constant.
Suppose that for $f\in\mathcal{F}$ with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\zeta$ there exists no function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$, then there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\alpha$ such that for each $S\in\mathcal{T}$ there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ and $z\in\mathcal{Y}$ such that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma\|p(\mathbf{x}^{U})\|_{2}$, where $U=(V+W)\cap V^{\perp}$.
@@
Since $\mathcal{F}$ satisfies Definition 3.2, with probability at least $\alpha$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, independent of $\mathbf{x}$, there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that
@@
In order to prove the statement, we need to show that the above condition can be extended from a single point to encompass an entire cube $S$. To achieve this, we introduce a random variable $y^{\prime}$, which is $(\zeta+O(\epsilon))$-close to $f$, such that, when conditioned on any $S\in\mathcal{S}$, $y^{\prime}$ is distributed uniformly within $S$ along the directions defined by $V$, while the other directions remain consistent with those of $y$.
@@
In the following claim we will prove that $f(\mathbf{x})$ and $f(\mathbf{x}^{\prime})$ are approximately close.
@@
It holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]=O(\epsilon)$.
@@
For each $i\in\mathcal{Y}$, define the boolean functions $f_{i}(\mathbf{x})\coloneqq\mathds{1}(f(\mathbf{x})=i)$ and let $g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})\coloneqq T_{\rho}f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})$, where $T_{\rho}$ is the Gaussian noise operator as given in Definition A.3. Fix a point $\mathbf{x}_{0}\in\mathbb{R}^{d}$ and a cube $S\in\mathcal{S}$. We first show that for any two points $\mathbf{x},\mathbf{z}\in S$, the difference $\lvert g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})-g_{\mathbf{x}_{0}}^{(i)}(\mathbf{z})\rvert$ is small.
By using that the gradient of $g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})$ is bounded by $1/\rho$ and that the distance between any $\mathbf{x},\mathbf{z}\in S$ is at most $\|\mathbf{x}-\mathbf{x}_{S}\|\leq\sqrt{k}\eta$, we get that
@@
for any $\mathbf{x},\mathbf{\mathbf{z}}\in S$. Consequently by the triangle inequality and Claim A.7, we have
$\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{E}}_{\mathbf{x}_{0},\mathbf{x}^{V}}[\lvert(f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})-g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x}))\mathds{1}(\mathbf{x}\in S)\rvert]$
@@
$\leq\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{E}}_{\mathbf{x}_{0},\mathbf{x}^{V}}[\lvert(f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})-g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x}_{S}))\mathds{1}(\mathbf{x}\in S)\rvert]+\eta\sqrt{k}/\rho\lesssim\sqrt{\rho}\Gamma(f_{i})+\eta\sqrt{k}/\rho\;,$
@@
for any collection of points $\mathbf{x}_{S}\in S$. Furthermore, by linearity of expectation and the fact that $\sum_{i\in\mathcal{Y}}f_{i}(\mathbf{x})=1$, we have that $\sum_{i\in\mathcal{Y}}g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})=1$ and $g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})\in[0,1]$ for all $\mathbf{x},\mathbf{x}_{0}\in\mathbb{R}^{d},i\in\mathcal{Y}$. Therefore, we can consider the multiclass function $G:\mathbb{R}^{d}\to\mathcal{Y}$ such that $G(\mathbf{x})$ is a random label sampled from the probability distribution induced by $g_{\mathbf{x}}^{(i)}(\mathbf{x}_{S})$, where $S\in\mathcal{S}$ such that $\mathbf{x}\in S$. Note that for every probability vector $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ and label $z\in\mathcal{Y}$ we have that $\|\mathbf{p}-\mathbf{e}_{z}\|_{1}=2\operatorname*{\mathbf{Pr}}_{i\sim\mathbf{p}}[i\neq z]$. Therefore, we have that the $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq G(\mathbf{x})]\lesssim\sqrt{\rho}\Gamma(f)+\eta\sqrt{k}\lvert\mathcal{Y}\rvert/\rho\lesssim\epsilon$.
@@
Notice that since $\mathbf{x}^{\prime}$ is distributed as $\mathcal{N}(\mathbf{0},\mathbf{I})$, we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x}^{\prime})]=O(\epsilon)$. However, as $G(\mathbf{x})$ is independent $\mathbf{x}^{V}$, we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x}^{\prime})]=\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x})]=O(\epsilon)$. Therefore, by the triangle inequality we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]=O(\epsilon)$, which completes the proof of Claim 3.6.∎
@@
Define $y^{\prime}$ such that or all $\mathbf{z}\in\mathbb{R}^{d}$, the distribution of $y^{\prime}$ given $\mathbf{x}^{\prime}=\mathbf{z}$ is the same as the distribution of $y$ given $\mathbf{x}=\mathbf{z}$. By the definition of $y^{\prime}$, it follows that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}^{\prime},y^{\prime}}[f(\mathbf{x}^{\prime})\neq y^{\prime}]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}^{\prime}}\left[\operatorname*{\mathbf{E}}_{y^{\prime}}[f(\mathbf{x}^{\prime})\neq y^{\prime}\mid\mathbf{x}^{\prime}]\right]\leq\zeta$. Therefore, it holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y^{\prime}]\leq\zeta+O(\epsilon)$.
@@
Therefore, with probability at least $\alpha$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ independent of $\mathbf{x}$, there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ and a label $y\in\mathcal{Y}$ such that
@@
However, since the moments of $y^{\prime}$ are identical when conditioned on each $\mathbf{x}_{0}^{V}$ that belongs to the same region in $S$, we deduce that if the moment condition holds for some polynomial at a particular point $\mathbf{x}_{0}^{V}$, then it must hold for the entire cube $S$, such that $\mathbf{x}_{0}^{V}\in S$. Therefore, there exists a subset $\mathcal{T}\subseteq\mathcal{S}$ of the partition with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\alpha$ satisfying the following: for all $S\in\mathcal{T}$ there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]>\sigma\|p(\mathbf{x}^{U})\|_{2}\;.$
@@
The remaining part of the section is structured as follows: In Section 3.1, we provide the analysis of Algorithm 4 with Proposition 3.7. In Section 3.2, we analyze Algorithm 3 providing the full proof of Theorem 3.3.